cindy had p100,less than Carl p250 more than Ronie.If the total amount of their money is p900, how much money had ronie?
Answer by algebra hello(55)

Let ronie had px,then according to question x-100 x 250 x=900 3x 150=900 3x=900-150 x=750/3 x=250 thus ronie had p250